a) y - y - 12y = 0 ; Y(0) = 3 and yo) = 5

b) y - 6y + 8y = 0 ; Y0 = 1 and yo) = 6

c) y - 12y + y = 0 ; YO) = 2 and y'_{(0)}=-3
Answer #1

@ you - 4* –12y =o , 7(0) = 3 and y(o)=5 cso Let y = emx be a solution of y = mene y = ma enne So from 0 (ma _m-12) ent=0 1Ⓡ y“ – 64+8y=0 , 7(0) = 1, Y CO)=6 In this case, we have m-6m +8=0 m2 - 4m -2m +8=0 = m (m-4) – a(m-4)=0 (m-2) (m-4)=0 1. Th1. © ya ley +9y = 0, y Co) =2, 8(0) = -30 The auxiliary equation in this case is m?–12m +9 = 0 . i m = 12 + 5144 – 36 127 11